In a moderate wage settlement, it was agreed that the wages of workers in the sector would be increased by 2.8% over the next four years. What percentage of wages increase on average per year?


1
Expert's answer
2021-06-08T14:49:26-0400

Let "p=" average annual percent of wagws increase

Then


"(1+\dfrac{p}{100})^4=(1+\dfrac{2.8}{100})"

"p=100(\sqrt[4]{1.028}-1)"

"p\approx0.6928\%"

Wages increase on average per year is "0.6928\ \%" per year.